Find the sum of the geometric series: 1+ sqrt6+ 6+...+7766
Answer by stanbon(75887)

Find the sum of the geometric series:
1+ sqrt6+ 6+...+7766
----
S(n) = a[(1-r^n)(1-r)]
---
Determine r and n
r = sqrt(6)/1 = sqrt(6)
n = ?
Solve (sqrt(6))^n = 7766 for "n":
n = log(7766)/(log(sqrt(6)) = 10
-----
S(11) = 1[(1-(sqrt(6))^11)/(1-sqrt(6)) = 19046.2322/1.4495 = 13,140
===================
Cheers,
Stan H.
================